Problem 17.5A Estimating inventory by the retail method. LO 17-5 The August 1 inventory of Hawkins Company had a cost of $81,Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Analyze Compute the retaiComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Analyze Compute the approComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Analyze Compute the costComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Analyze What is the amoun
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Hawkins
Cost Retail Note
Beginning Inventory      81,000.00 116,000.00
Plus: Purchases      97,710.00 143,000.00
Goods available for sale 178,710.00 259,000.00 A
Cost to retail % 69.00% B This is Goods available for sale- Cost/ Goods available for sale- Retail i.e. $ 178,710/ $ 259,000.
Less: Net sales 127,000.00 F
Estimated Ending Inventory- Retail 132,000.00 C
Estimated Ending Inventory- Cost      91,080.00 D= C* B
Estimated Cost of goods sold     87,630.00 E=A-D
Estimated Gross profit     39,370.00 G=F-E